For a single phase transformer with 250 primary turns and 50 secondary turns connected across a 1500V,50Hz supply, the maximum value of flux is


1
Expert's answer
2021-03-18T03:37:33-0400

"E=N(d\Phi/dt)=N\omega\Phi_m\cos(\omega t)"


"E_m=N\omega\Phi_m"


"E_{rms}=\frac{N\omega\Phi_m}{\sqrt2}=\frac{N2\pi f\Phi_m}{\sqrt2}=4.44Nf\Phi_m\to"


"\Phi_m=\frac{E_{rms}}{4.44Nf}=\frac{1500}{4.44\cdot 250\cdot 50}=0.027\ (Wb)" . Answer